# Under the null of valid specification, pre-tests of valid specification do not distort inference

### v1

@@ 2 The set-up and main result
Suppose that Assumption 1 holds and $P_{U}\in\mathcal{P}_{0}$. Then, for any convex set $C$ that is symmetric around the origin,
$\lim_{n}P\left[\sqrt{n}\big{(}\widehat{\beta}-\beta_{0}\big{)}\in C|T_{1,n}\leq q_{1,n},...,T_{J,n}\leq q_{J,n}\right]\geq P(Z\in C),$
where $Z\sim\mathcal{N}(0,V(\psi))$. [^3] Assumption 1 does not guarantee that the test statistics $T_{j,n}$ are measurable. If not, the result holds replacing $P$ by the outer probability $P^{*}$ (see, e.g. Van der Vaart and Wellner, 2023, p.6).
The result relies on two main ingredients. To better understand them, suppose for simplicity that $J=1$. The first main ingredient is the approximation of the limit distribution of $T_{1}(n^{1/2}P_{n})$ by a convex function of a finite-dimensional Gaussian measure. The second ingredient is the Gaussian correlation inequality (Royen, 2014), which states that for any centered Gaussian distribution $\mu$ on $\mathbb{R}^{q}$ and two convex sets $E$ and $F$ that are symmmetric about the origin,
@@
The main implication of Theorem 3 concerns (conditional) inference on $\beta_{0}$. Let $\widehat{V}$ denote an estimator $V(\psi(U_{1}))$. We consider the usual F-test of $\beta_{0}=b_{0}$ with test statistic $F_{n}(b_{0}):=(\widehat{\beta}-b_{0})^{\prime}\widehat{V}^{-1}(\widehat{\beta}-b_{0})$ and critical region $\{F_{n}(b_{0})>q_{1-\alpha}(p)\}$, with $q_{1-\alpha}(p)$ the quantile of order $1-\alpha$ of a chi-squared distribution with $p$ degrees of freedom. We also consider the standard confidence region
$\text{CR}_{1-\alpha}=\left\{b\in\mathbb{R}^{p}:n(\widehat{\beta}-b)^{\prime}\widehat{V}^{-1}(\widehat{\beta}-b)\leq q_{1-\alpha}(p)\right\}.$
@@
Suppose that Assumption 1 holds, $P_{U}\in\mathcal{P}_{0}$ and $\widehat{V}\stackrel{{\scriptstyle p}}{{\longrightarrow}}V(\psi(U_{1}))$. Then:
@@
Hence, under Assumption 1 and provided that $P_{U}\in\mathcal{P}_{0}$, conditional on accepting specification tests the usual $F$-tests and confidence regions are asymptotically conservative.
3 Extensions
While quite general, the previous set-up excludes some important cases, that we cover now under the following alternate assumption.
@@ 3 Extensions
    We observe a sample $(U_{i})_{i=1,...,n}$ of identically distributed random vectors with probability distribution $P_{U}$.
@@
    If $P_{U}\in\mathcal{P}_{0}$, we have,
    $\left(\widehat{\Sigma}_{\beta}^{-1/2}\left(\widehat{\beta}-\beta_{0}\right),\widehat{\Sigma}_{\theta}^{-1/2}\left(\widehat{\theta}-\theta_{0}\right)\right)\stackrel{{\scriptstyle d}}{{\longrightarrow}}\mathcal{N}\left(0,\begin{pmatrix}I_{p}&\Sigma_{12}\\ \Sigma_{12}&I_{q}\end{pmatrix}\right),$
    where $I_{k}$ is the identity matrix of size $k$.
@@
    If $P_{U}\in\mathcal{P}_{1}$, with $\mathcal{P}_{0}\subseteq\mathcal{P}_{1}$, we have $\theta_{0}=0$.
@@
    We consider $J$ tests of $H_{0}:\;P_{U}\in\mathcal{P}_{1}$ based on the statistics $(T_{1,n},...,T_{J,n})\in\mathbb{R}^{J}$ that satisfy, under $H_{0}$,
    $T_{j,n}=T_{j}\left(\widehat{\Sigma}_{\theta}^{-1/2}\left(\widehat{\theta}-\theta_{0}\right)\right)+o_{P}(1),$ (3)
    where $P_{n}$ denotes the empirical measure of $U$, indexed by $\mathcal{F}$. Moreover, $T_{j}$ is convex on $\mathbb{R}^{q}$ and satisfies $T_{j}(-x)=T_{j}(x)\geq 0$ for all $x\in\mathbb{R}^{q}$ and $T_{j}(0)=0$.
  5. 5.
    The critical region of the $j$-th test of $H_{0}$ is $\{T_{j,n}>q_{j,n}\}$, where $q_{j,n}$ is a random variable satisfying $q_{j,n}\stackrel{{\scriptstyle p}}{{\longrightarrow}}q_{j}>0$.
Unlike Assumption 1, Assumption 2 allows for estimators that converge at rates lower than $n^{1/2}$. An important example is regression discontinuity designs, where specification tests would either be tests that the density is continuous at the threshold (McCrary, 2008) or tests that predetermined variables have continuous means at the threshold. Assumption 2 allows for this case, when the testable implications are of finite dimension, denoted by $q$. Assumption 2 also allows for non-independent but identically distributed observations, as is for instance the case with stationary time series, provided we can prove the asymptotic normality of the actual and specification-test estimators.
@@
Suppose that Assumption 2 holds and $P_{U}\in\mathcal{P}_{0}$. Then, for any convex set $C$ that is symmetric around the origin,
$\lim_{n}P\left[\widehat{\Sigma}_{\beta}^{-1/2}\left(\widehat{\beta}-\beta_{0}\right)\in C|T_{1,n}\leq q_{1,n},...,T_{J,n}\leq q_{J,n}\right]\geq P(Z\in C),$
where $Z\sim\mathcal{N}(0,I_{p})$.
As above, this theorem implies conservative tests and confidence regions on $\beta_{0}$ conditional on not rejecting specification tests.
Appendix A Proofs
A.1 Proof of Theorem 3
Let us define $Z_{n}:=\sqrt{n}(\widehat{\beta}-\beta_{0})$, $T_{n}:=\max_{j=1,...,J}[T_{j,n}-q_{j,n}]$ and $T(P):=\max_{j=1,...,J}[T_{j}(P)-q_{j}]$. $T$ is convex and continuous, as the maximum of $J$ convex and continuous functionals. Moreover,
@@ A.1 Proof of Theorem 3
$\partial(C\times(-\infty,0])=(\partial C\times(-\infty,0])\cup(C\times\{0\}).$
Moreover, since $C$ is convex, $\partial C$ has Lebesgue measure zero (see, e.g. Lang, 1986, p.90). Since $V(\psi(U_{1}))$ is nonsingular, the distribution of $G\psi$ is absolutely continuous with respect to the Lebesgue measure. Hence, $P(G\psi\in\partial C)=0$. By Theorem 11.1 and Problem 11.3 in Davydov et al. (1998) and continuity of $T$, the cumulative distribution function $H$ of $T(G)$ is strictly increasing on $(r_{0},\infty)$, with $r_{0}:=\inf_{P}T(P)=\max_{j=1,...,J}-q_{j}<0$. Hence, $H$ is continuous at 0, $H(0)>0$ and $P(T(G)=q)=0$. Hence,
@@
$\lim_{n}P^{*}(Z_{n}\in C,T_{n}\leq 0)=P(G\psi\in C,T(G)\leq 0),$
where $P^{*}$ denotes the outer probability. Similarly, $\lim_{n}P^{*}(T_{n}\leq 0)=P(T(G)\leq 0)=H(0)>0$. Hence,
$\lim_{n}P^{*}(Z_{n}\in C|T_{n}\leq 0)=P(G\psi\in C|T(G)\leq 0).$ (5)
@@
$=$ $\,P(Z\in C)P(T(\widetilde{G})\leq-\varepsilon),$
where the last equality follows by definition of $Z$. Hence,
$P(G\psi\in C|\;T(G)\leq 0)\geq$ $P(Z\in C)\frac{P(T(\widetilde{G})\leq-\varepsilon)}{H(0)}$
$\geq$ $P(Z\in C)\frac{P(T(G)\leq-2\varepsilon)}{H(0)}$
$\geq$ $P(Z\in C)\frac{H(-2\varepsilon)}{H(0)}.$
@@
A.2 Proof of Corollary 1
Let =$\mathcal{E}=\left\{x:\,x^{\prime}V^{-1}x\leq q_{1-\alpha}(p)\right\}$. By assumption, under the null hypothesis $b_{0}=\beta_{0}$,
$F_{n}(b_{0})=n(\widehat{\beta}-\beta_{0})^{\prime}V^{-1}(\widehat{\beta}-\beta_{0})+o_{P}(1).$
As a result,
$P\left(F_{n}(b_{0})\leq q_{1-\alpha}(p)|T_{1,n}\leq q_{1,n},...,T_{J,n}\leq q_{J,n}\right)-P\left(\sqrt{n}(\widehat{\beta}-\beta_{0})\in\mathcal{E}|T_{1,n}\leq q_{1,n},...,T_{J,n}\leq q_{J,n}\right)\to 0$
The result follows by Theorem 3.
A.3 Proof of Theorem 2
The proof is almost the same as that of Theorem 3, and we just highlight the differences. First, since $T_{j}$ is a convex function on $\mathbb{R}^{q}$, it is continuous. Hence, $T(x):=\max_{j=1,...,J}[T_{j}(x)-q_{j}]$ is continuous (and convex). Then, reasoning as above, we obtain
$\lim_{n\to\infty}P\left[\widehat{\Sigma}_{\beta}^{-1/2}\left(\widehat{\beta}-\beta_{0}\right)\in C|T_{n}\leq 0\right]=P(Z\in C|T(Z^{\prime})\leq 0),$
@@ A.3 Proof of Theorem 2
$(Z,Z^{\prime})\sim\mathcal{N}\left(0,\begin{pmatrix}I_{p}&\Sigma_{12}\\ \Sigma_{12}&I_{q}\end{pmatrix}\right).$
The result follows by the Gaussian correlation inequality.

### v2

@@ 2 Set-up and examples
which implies that $T_{1}$ is continuous.
Specification Test 5 (Kolmogorov-Smirnov test statistics with nuisance parameters)
In some cases, the restrictions $Pf=0$ hold for $f\in\mathcal{F}_{\eta_{0}}$, where $\eta_{0}$ is unknown but can be consistently estimated by $\widehat{\eta}$. In Example 5, for instance, $\eta_{0}=(\alpha_{0},\beta_{0})$, $\mathcal{F}_{\eta}=\{f_{d^{\prime},\eta}:(y,d)\mapsto\mathds{1}\left\{d\leq d^{\prime}\right\}(y-\eta_{1}-\eta_{2}d),\;d^{\prime}\in\mathcal{D}\}$ and $\widehat{\eta}$ is the OLS estimator in the regression of $Y$ on $X:=(1,D)^{\prime}$. In such cases, Assumption 2 can still hold under regularity conditions, provided that we replace $\mathcal{F}_{\eta_{0}}$ by another appropriate class. To see this, consider Example 5 again, with a Kolmogorov-Smirnov test statistic (the same reasoning applies to a Cramér-von Mises test statistic). Then $T_{1,n}=n^{1/2}\sup_{d\in\mathcal{D}}|\widehat{\theta}(d)|$, with
$\widehat{\theta}(d):=\frac{1}{n}\sum_{i=1}^{n}\mathds{1}\left\{D_{i}\leq d\right\}(Y_{i}-X_{i}^{\prime}\widehat{\eta}).$
By a uniform law of large numbers on $\sum_{i=1}\mathds{1}\left\{D_{i}\leq d\right\}X_{i}/n$ and standard results on regressions, we obtain
$\sqrt{n}\left(\widehat{\theta}(d)-\theta_{0}(d)\right):=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\left[\mathds{1}\left\{D_{i}\leq d\right\}+E[\mathds{1}\left\{D\leq d\right\}X^{\prime}]E[XX^{\prime}]^{-1}X_{i}\right](Y_{i}-X_{i}^{\prime}\eta_{0})+o_{P}(1),$
where the $o_{P}(1)$ is uniform over $d\in\mathcal{D}$. Let us define
$\widetilde{\mathcal{F}}:=$$\bigg{\{}\widetilde{f}_{d^{\prime}}:(y,d)\mapsto\left[\mathds{1}\left\{d\leq d^{\prime}\right\}+E[\mathds{1}\left\{D\leq d^{\prime}\right\}X^{\prime}]E[XX^{\prime}]^{-1}(1,d)^{\prime}\right](y-\alpha_{0}-\beta_{0}d),$
$\quad d^{\prime}\in\mathcal{D}\bigg{\}}.$
Note that we still have $Pf=0$ for all $f\in\widetilde{\mathcal{F}}$, and like $\mathcal{F}_{\eta}$, $\widetilde{\mathcal{F}}$ is Donsker. Moreover, by continuity of the supremum functional, $T_{1,n}=T_{1}(P)+o_{P}(1)$, with $T_{1}(P)=\sup_{f\in\widetilde{\mathcal{F}}}|Pf|$. Hence, Point 3 of Assumption 2 and the restrictions on $T_{1,n}$ in Point 4 of that assumption hold with $\mathcal{F}=\widetilde{\mathcal{F}}$. To obtain similar results in other, potentially more complicated setups, see for instance Chapter 13 in Van der Vaart and Wellner (2023).
3 Results
Our first result gives a control on the asymptotic probability that the properly normalized estimator of interest belongs to a convex set, conditional on not rejecting the specification tests.
@@ 3 Results
Suppose that $P_{U}\in\mathcal{P}_{0}$. Then, for any convex set $C$ that is symmetric around the origin,
  1. 1.
    If Assumption 1 holds,
    $\lim_{n\to\infty}P\left[\widehat{\Sigma}_{\beta}^{-1/2}\left(\widehat{\beta}-\beta_{0}\right)\in C\big{|}T_{1,n}\leq q_{1,n},...,T_{J,n}\leq q_{J,n}\right]\geq P(Z_{1}\in C),$
    where $Z_{1}\sim\mathcal{N}(0,I_{p})$.
  2. 2.
    If Assumption 2 holds,
    $\lim_{n\to\infty}P\left[\sqrt{n}\big{(}\widehat{\beta}-\beta_{0}\big{)}\in C\big{|}T_{1,n}\leq q_{1,n},...,T_{J,n}\leq q_{J,n}\right]\geq P(Z_{2}\in C),$ (6)
    where $Z_{2}\sim\mathcal{N}(0,V(\psi))$. [^6] Assumption 2 does not guarantee that the test statistics $T_{j,n}$ are measurable. If not, the result holds replacing $P$ on the left-hand side of (6) by the outer probability $P^{*}$ (see, e.g. Van der Vaart and Wellner, 2023, p.6). The same remark applies to Corollary 1 below.
The first point follows in particular from properties of convex functionals of Gaussian measures (Davydov et al., 1998), and the Gaussian correlation inequality (Royen, 2014). The latter states that for any centered Gaussian distribution $\mu$ on $\mathbb{R}^{q}$ and two convex sets $E$ and $F$ that are symmetric about the origin,
@@
The second result follows from the same tools, but also from the approximation of the limit distribution of $T_{j}(n^{1/2}P_{n})$ by a convex function of a finite-dimensional Gaussian measure.
We now turn to the implications of Theorem 1 for (conditional) inference on $\beta_{0}$. Let $\widehat{V}:=\widehat{\Sigma}_{\beta}$ if Assumption 1 holds; otherwise, let $\widehat{V}$ denote an estimator of $V(\psi(U_{1}))/n$. We consider the usual F-test of $\beta_{0}=b_{0}$ with test statistic $F_{n}(b_{0}):=(\widehat{\beta}-b_{0})^{\prime}\widehat{V}^{-1}(\widehat{\beta}-b_{0})$ and critical region $\{F_{n}(b_{0})>q_{1-\alpha}(p)\}$, with $q_{1-\alpha}(p)$ the quantile of order $1-\alpha$ of a chi-squared distribution with $p$ degrees of freedom. We also consider the standard confidence region
$\text{CR}_{1-\alpha}=\left\{b\in\mathbb{R}^{p}:(\widehat{\beta}-b)^{\prime}\widehat{V}^{-1}(\widehat{\beta}-b)\leq q_{1-\alpha}(p)\right\}.$
@@
Suppose that either Assumption 1 or Assumption 2 holds, $P_{U}\in\mathcal{P}_{0}$ and, if Assumption 2 holds, $n\widehat{V}\stackrel{{\scriptstyle P}}{{\longrightarrow}}V(\psi(U_{1}))$. Then:
@@
Hence, under Assumption 2 and provided that $P_{U}\in\mathcal{P}_{0}$, conditional on accepting specification tests the usual $F$-tests and confidence regions are asymptotically conservative. In the common case of inference on a single coefficient ($p=1$), this implies that two-sided tests and confidence intervals on $\beta_{0}$ are asymptotically conservative.
On the other hand, Corollary 1 does not provide guarantees for one-sided tests and confidence intervals. It turns out, however, that we can obtain such guarantees by the following consequence of the Gaussian correlation inequality. [^7] We thank Fedor Petrov for giving us the proof of this result.
Lemma 1
Let $\mu$ denote a mean zero Gaussian measure on $\mathbb{R}^{q+1}$, $K=(-\infty,a]\times\mathbb{R}^{q}$ with $a\geq 0$ and $L=\mathbb{R}\times C$ where $C\subset\mathbb{R}^{q}$ is convex and symmetric around the origin. Then $\mu(K\cap L)\geq\mu(K)\mu(L)$. The same result holds with $K=[-a,\infty)\times\mathbb{R}^{q}$.
In this unidimensional case ($p=1$), we define the test statistic $T_{n}(b_{0}):=(\widehat{\beta}-b_{0})/\widehat{V}^{1/2}$ and the critical region $\{T_{n}(b_{0})>z_{1-\alpha}\}$, with $z_{1-\alpha}$ the quantile of order $1-\alpha$ of a standard normal distribution (of course, the same result would hold with instead the critical region $\{T_{n}(b_{0})<z_{\alpha}\}$). Accordingly, we consider the unilateral confidence interval
$\text{CI}_{1-\alpha}=(-\infty,\widehat{\beta}+\widehat{V}^{1/2}z_{1-\alpha}].$
Proposition 1
Suppose that $\alpha\in(0,1/2]$, $p=1$ and the assumptions in Corollary 1 hold. Then:
  1. 1.
    If $\beta_{0}=b_{0}$, $\lim_{n\to\infty}P\left(T_{n}(b_{0})>z_{1-\alpha}(p)|T_{1,n}\leq q_{1,n},...,T_{J,n}\leq q_{J,n}\right)\leq\alpha$.
  2. 2.
    $\lim_{n\to\infty}P\left(\beta_{0}\in\text{CI}_{1-\alpha}|T_{1,n}\leq q_{1,n},...,T_{J,n}\leq q_{J,n}\right)\geq 1-\alpha$.
So far, we have considered pointwise results, where the distribution $P_{U}$ does not depend on $n$, the sample size. In Appendix A, we consider a setup where $P_{U}$ may vary with the sample size. We show, under a suitable adaptation of Assumption 1, that the same results as above hold, but now uniformly over $\mathcal{P}_{0}$.
Appendix A Uniform result
We consider here a setting where the probability distribution of the data changes with the sample size; accordingly, we denote it by $P_{U_{n}}$ instead of $P_{U}$. In this setup, Assumption 3 is the exact counterpart of Assumption 1. Note that we emphasize the dependence of the true parameters in $P_{U_{n}}$ by denoting them now by $\beta(P_{U_{n}})$ and $\theta(P_{U_{n}})$.
@@ Appendix A Uniform result
    For each $n\geq 1$, we observe a sample $(U_{n,i})_{i=1,...,n}$ of identically distributed random vectors ($U_{n,i}\in\mathbb{R}^{k}$) with probability distribution $P_{U_{n}}(A):=P(U_{n,1}\in A)$ for all Borel set $A\subset\mathbb{R}^{k}$.
@@
    For any sequence $(P_{U_{n}})_{n\geq 1}$ with $P_{U_{n}}\in\mathcal{P}_{0}$ for all $n\geq 1$, we have
    $\left(\widehat{\Sigma}_{\beta}^{-1/2}\left(\widehat{\beta}-\beta(P_{U_{n}})\right),\widehat{\Sigma}_{\theta}^{-1/2}\left(\widehat{\theta}-\theta(P_{U_{n}})\right)\right)\stackrel{{\scriptstyle d}}{{\longrightarrow}}\mathcal{N}\left(0,\begin{pmatrix}I_{p}&\Sigma_{12}\\ \Sigma_{12}^{\prime}&\Sigma_{22}\end{pmatrix}\right),$
    where we use the same notation on matrices as in Assumption 1.
@@
    If $P_{U_{n}}\in\mathcal{P}_{0}$, we have $\theta(P_{U_{n}})=0$.
@@
    We consider $J$ specification tests of $H_{0}:P_{U_{n}}\in\mathcal{P}_{0}$ based on the statistics $(T_{1,n},...,T_{J,n})\in\mathbb{R}^{J}$ that satisfy, under $H_{0}$,
    $T_{j,n}=T_{j}\left(\widehat{\Sigma}_{\theta}^{-1/2}\widehat{\theta}\right)+o_{P}(1),$
    and $T_{j}$ satisfies the same restrictions as in Assumption 1. The critical region of the $j$-th test is $\{T_{j,n}>q_{j,n}\}$, where the random variables $q_{j,n}$ satisfy $q_{j,n}\stackrel{{\scriptstyle P}}{{\longrightarrow}}q_{j}>0$.
@@
Suppose that Assumption 3 holds and $P_{U_{n}}\in\mathcal{P}_{0}$ for all $n\geq 1$. Then, for any convex set $C$ that is symmetric around the origin,
$\lim_{n\to\infty}P\left[\widehat{\Sigma}_{\beta}^{-1/2}\left(\widehat{\beta}-\beta(P_{U_{n}})\right)\in C\big{|}T_{1,n}\leq q_{1,n},...,T_{J,n}\leq q_{J,n}\right]\geq P(Z_{1}\in C),$
where $Z_{1}\sim\mathcal{N}(0,I_{p})$. Moreover,
$\limsup_{n\to\infty}\sup_{P_{U_{n}}\in\mathcal{P}_{0}}$ $P\left(F_{n}(\beta(P_{U_{n}}))>q_{1-\alpha}(p)|T_{1,n}\leq q_{1,n},...,T_{J,n}\leq q_{J,n}\right)\leq\alpha,$ (7)
$\liminf_{n\to\infty}\inf_{P_{U_{n}}\in\mathcal{P}_{0}}$ $P\left(\beta(P_{U_{n}}))\in\text{CR}_{1-\alpha}|T_{1,n}\leq q_{1,n},...,T_{J,n}\leq q_{J,n}\right)\geq 1-\alpha.$ (8)
Finally, if $p=1$ (7) (resp. (8)) also holds true with $T_{n}(\beta(P_{U_{n}}))$ instead of $F_{n}(\beta(P_{U_{n}}))$ (resp. $\text{CI}_{1-\alpha}$ instead of $\text{CR}_{1-\alpha}$).
The first part of Theorem 2 establishes that Point 1 of Theorem 1 holds along any sequence $(P_{U_{n}})_{n\geq 1}$ so long as $P_{U_{n}}\in\mathcal{P}_{0}$. The second and third results are similar to those of Corollary 1, but now hold uniformly over $\mathcal{P}_{0}$.
Appendix B Proofs
B.1 Theorem 1
Point 2
We begin by the proof of Point 2 as this is more complicated case. Let us define $Z_{n}:=\sqrt{n}(\widehat{\beta}-\beta_{0})$, $T_{n}:=\max_{j=1,...,J}[T_{j,n}-q_{j,n}]$ and $T(P):=\max_{j=1,...,J}[T_{j}(P)-q_{j}]$. $T$ is convex and continuous, as the maximum of $J$ convex and continuous functionals. Moreover,
@@ B.1 Theorem 1
$\partial(C\times(-\infty,0])=(\partial C\times(-\infty,0])\cup(C\times\{0\}).$ (10)
Moreover, since $C$ is convex, $\partial C$ has Lebesgue measure zero (see, e.g. Lang, 1986, p.90). Since $V(\psi(U_{1}))$ is nonsingular, the distribution of $G\psi$ is absolutely continuous with respect to the Lebesgue measure. Hence, $P(G\psi\in\partial C)=0$. By Theorem 11.1 and Problem 11.3 in Davydov et al. (1998) and continuity of $T$, the cumulative distribution function $H$ of $T(G)$ is strictly increasing on $(r_{0},\infty)$, with $r_{0}:=\inf_{P\in\ell^{\infty}(\mathcal{F})}T(P)=\max_{j=1,...,J}-q_{j}<0$. Hence, $H$ is continuous at 0, $H(0)>0$ and $P(T(G)=0)=0$. Hence, in view of (10),
@@
$\lim_{n\to\infty}P^{*}(Z_{n}\in C,T_{n}\leq 0)=P(G\psi\in C,T(G)\leq 0),$
where $P^{*}$ denotes the outer probability. Similarly, $\lim_{n\to\infty}P^{*}(T_{n}\leq 0)=P(T(G)\leq 0)=H(0)>0$. Hence,
$\lim_{n\to\infty}P^{*}(Z_{n}\in C|T_{n}\leq 0)=P(G\psi\in C|T(G)\leq 0).$ (11)
@@
$=$ $\,P(Z_{2}\in C)P(T(\widetilde{G})\leq-\varepsilon),$
where the last equality follows by definition of $Z_{2}$. Hence,
$P(G\psi\in C\,|\,T(G)\leq 0)\geq$ $P(Z_{2}\in C)\frac{P(T(\widetilde{G})\leq-\varepsilon)}{H(0)}$
$\geq$ $P(Z_{2}\in C)\frac{P(T(G)\leq-2\varepsilon)}{H(0)}$
$\geq$ $P(Z_{2}\in C)\frac{H(-2\varepsilon)}{H(0)}.$
@@
Point 1
The begining of the proof is almost the same as that in Point 2, so we just highlight the differences. First, since $T_{j}$ is a convex function on $\mathbb{R}^{q}$, it is continuous. Hence, $T(x):=\max_{j=1,...,J}[T_{j}(x)-q_{j}]$ is continuous (and convex). Then, reasoning as above, we obtain
$\lim_{n\to\infty}P\left[\widehat{\Sigma}_{\beta}^{-1/2}\left(\widehat{\beta}-\beta_{0}\right)\in C\,\big{|}\,T_{n}\leq 0\right]=P(Z_{1}\in C\,|\,T(Z_{1}^{\prime})\leq 0),$
@@
$(Z_{1},Z_{1}^{\prime})\sim\mathcal{N}\left(0,\begin{pmatrix}I_{p}&\Sigma_{12}\\ \Sigma_{12}^{\prime}&\Sigma_{22}\end{pmatrix}\right).$
The result follows directly by the Gaussian correlation inequality.
B.2 Proof of Corollary 1
Point 2 follows directly from Point 1 so we focus on the latter. Suppose first that Assumption 1 holds. Let $C=\{x\in\mathbb{R}^{p}:x^{\prime}x\leq q_{1-\alpha}(p)\}$. Then, if $\beta_{0}=b_{0}$,
$F_{n}(b_{0})\leq q_{1-\alpha}(p)\Leftrightarrow\widehat{\Sigma}_{\beta}^{-1/2}\left(\widehat{\beta}-\beta_{0}\right)\in C.$
Because $C$ is convex and symmetric around the origin, we have, by Theorem 1,
$\lim_{n\to\infty}P\left[F_{n}(b_{0})\leq q_{1-\alpha}(p)\big{|}T_{1,n}\leq q_{1,n},...,T_{J,n}\leq q_{J,n}\right]$
$=$ $\lim_{n\to\infty}P\left[\widehat{\Sigma}_{\beta}^{-1/2}\left(\widehat{\beta}-\beta_{0}\right)\in C\,\big{|}\,T_{1,n}\leq q_{1,n},...,T_{J,n}\leq q_{J,n}\right]$
$\geq$ $P(Z_{1}\in C)=1-\alpha.$
The result follows in this case.
Now, suppose that Assumption 2 holds. Because $n\widehat{V}\stackrel{{\scriptstyle P}}{{\longrightarrow}}V(\psi(U_{1}))$ and $V(\psi(U_{1}))$ is nonsingular, we have, if $\beta_{0}=b_{0}$,
$F_{n}(b_{0})=n(\widehat{\beta}-\beta_{0})^{\prime}V(\psi(U_{1}))^{-1}(\widehat{\beta}-\beta_{0})+o_{P}(1).$
Let $C=\left\{x:\,x^{\prime}V(\psi(U_{1}))^{-1}x\leq q_{1-\alpha}(p)\right\}$. $C$ is convex and symmetric around the origin. Then, by the previous display and Theorem 1,
$\lim_{n\to\infty}P\left(F_{n}(b_{0})\leq q_{1-\alpha}(p)\,\big{|}\,T_{1,n}\leq q_{1,n},...,T_{J,n}\leq q_{J,n}\right)$
$=$ $\lim_{n\to\infty}P\left(\sqrt{n}(\widehat{\beta}-\beta_{0})\in C\,\big{|}\,T_{1,n}\leq q_{1,n},...,T_{J,n}\leq q_{J,n}\right)$
$\geq$ $P(Z_{2}\in C)=1-\alpha.$
The result follows.
B.3 Proof of Lemma 1
Let $K_{0}=[-a,a]\times\mathbb{R}^{k-1}$, $K_{-}=(-\infty,-a]\times\mathbb{R}^{k-1}$ and $K_{+}=[a,\infty)\times\mathbb{R}^{k-1}$. By the Gaussian correlation inequality and symmetry of $\mu$,
$\mu(L)\left[\mu(K_{0})+2\mu(K_{+})\right]$ $=\mu(L)$
$=\mu(L\cap K_{0})+\mu(L\cap K_{-})+\mu(L\cap K_{+})$
$\geq\mu(L)\mu(K_{0})+2\mu(L\cap K_{+}).$
Hence, $\mu(L)\mu(K_{+})\geq\mu(L\cap K_{+})$. As a result,
$\mu(L\cap K)$ $=\mu(L)-\mu(L\cap K_{+})$
$\geq\mu(L)-\mu(L)\mu(K_{+})$
$=\mu(L)\mu(K).$
B.4 Proof of Proposition 1
By applying Lemma 1 instead of the initial Gaussian correlation inequality, we obtain the same result as Theorem 1 but with $C=(-\infty,a]$. Then, the same reasoning as in Corollary 1 leads to the result.
B.5 Proof of Theorem 2
The proof the first result is exactly the same as that of Point 1 of Theorem 1. To show (7), fix $\eta>0$ and pick a sequence $(Q_{U_{n}})_{n\geq 1}$ in $\mathcal{P}_{0}$ such that for all $n\geq 1$,
$P\left(F_{n}(\beta(Q_{U_{n}}))>q_{1-\alpha}(p)|T_{1,n}\leq q_{1,n},...,T_{J,n}\leq q_{J,n}\right)$
$\geq$ $\sup_{P_{U_{n}}\in\mathcal{P}_{0}}P\left(F_{n}(\beta(P_{U_{n}}))>q_{1-\alpha}(p)|T_{1,n}\leq q_{1,n},...,T_{J,n}\leq q_{J,n}\right)-\eta.$ (12)
Applying the same reasoning as that used to prove Point 1 of Corollary 1, we obtain
$\lim_{n\to\infty}P\left(F_{n}(\beta(Q_{U_{n}}))>q_{1-\alpha}(p)|T_{1,n}\leq q_{1,n},...,T_{J,n}\leq q_{J,n}\right)\leq\alpha.$
Combined with (12), this yields
$\limsup_{n\to\infty}\sup_{P_{U_{n}}\in\mathcal{P}_{0}}P\left(F_{n}(\beta(P_{U_{n}}))>q_{1-\alpha}(p)|T_{1,n}\leq q_{1,n},...,T_{J,n}\leq q_{J,n}\right)\leq\alpha+\eta.$
Equation (7) follows since $\eta$ was arbitrary. The adaptation to $T_{n}(\beta(P_{U_{n}}))$, Equation (8) and its adaptation to $\text{CI}_{1-\alpha}$ instead of $\text{CR}_{1-\alpha}$ follow with the same reasoning.